OpenStudy (anonymous):

Solve using perfect square factoring patterns. y^2 + 16y + 64 = 0

OpenStudy (anonymous):

the perfect square pattern that can be used here is (a+b)^2=a2+2ab+b2.

OpenStudy (anonymous):

In the expression it will be: y2+16y+64=y2+2*8y+8^2=(y+8)^2

OpenStudy (anonymous):

the solution will be the value of x when (x+8) turns into 0, so the x=-8